# Super-resolving frequency measurement with mode-selective quantum memory

### v1

@@ II.2 Mode-selective Raman quantum memory
In the low-coupling regime (i.e. at low storage efficiencies), the memory operates in a single-mode fashion: the temporal mode of the signal stored in the memory is directly determined by the temporal mode of the control pulse [56] (see the Supplementary Materials for more discussions). As illustrated in Fig.1(b), when both the signal and control fields are in the HG0 mode (labeled as 1), the signal is stored and subsequently retrieved. In contrast, when the signal is in the HG0 mode but the control is in the HG1 mode (labeled as 2), there is nearly no storage or retrieval. Experimental results for these two scenarios are presented in Fig.1(c). To quantify the mode selectivity, we measured the total efficiencies for all combinations of signal and control pulses prepared in HG0 and HG1 modes. The efficiencies were normalized for each signal mode, producing the mode crosstalk matrix shown in Fig. 1(d). Our results confirm high mode selectivity: we achieved $99.61\pm 0.48\%$ mode selectivity for the storage of HG0 signal with HG0 control, and a selectivity of $0.39\pm 0.02\%$ when stored with HG1 control. These results were obtained using both read-in and read-out control pulses with energies of $125\pm 1$ pJ, yielding total efficiencies of approximately 11% along the diagonal terms. It is important to note that at higher coupling strengths (e.g., with stronger control pulses), this single-mode behavior degrades due to the increased storage of temporal modes orthogonal to that of the intended control field. This highlights an inherent trade-off between mode selectivity and overall memory efficiency. By achieving low crosstalk and high mode selectivity in the low-coupling regime, our Raman memory provides a robust and effective platform for the super-resolution task.
@@ II.3 Experimental superresolution with mode-selective quantum memory
Figure 2: Experimental estimation results. This graph compares the maximum likelihood estimators (red markers) with raw estimators (yellow markers), both derived from $10\times 10^{3}$ total detected photons. Error bars represent one standard deviation from 50 bootstrapped samples. The dashed black line indicates the ground truth separation. The shaded regions illustrate the theoretically predicted standard deviations from the quantum CRLB (blue) and the conventional direct intensity detection CRLB (gray).
@@
To perform estimation experiments using the mode-selective Raman quantum memory described above, we first prepared the signal as an incoherent mixture of two Gaussian spectral lines, each with a linewidth of $\sigma=5.30$ MHz (see Methods for detailed procedure). We varied their frequency separation across 21 values, ranging from 0 to 1 in steps of 0.05. For each signal, we performed two measurements, storing the signal with an HG0 or an HG1 control read-in pulse. In both cases, retrieval was conducted using a $\text{HG}_{1}$ control pulse after 200 ns storage and the retrieved signal was detected with superconducting nanowire single-photon detectors. At $\epsilon=0$, the storage efficiency and total internal efficiency for the $\text{HG}_{0}$ control read-in were measured to be $39.5\pm 0.6\%$ and $14.6\pm 0.1\%$ respectively, using control pulses with energies of $130\pm 1$ pJ. For each separation, we recorded approximately $N\approx 2\times 10^{5}$ detector clicks in total. To estimate the frequency separation $\epsilon$, we first performed a full calibration of the experimental system with $1.6\times 10^{5}$ detected counts for each separation. By fitting the experimental model to this calibration data, we determined the measured crosstalk from the $\text{HG}_{0}$ to $\text{HG}_{1}$ modes in the estimation experiment to be 0.34%. It is noteworthy that this is a separate experiment from the demonstration of mode filtering described in the previous section, accounting for the slight difference in the measured mode crosstalk values. With the calibrated model, in each run we performed MLE to compute the estimate $\hat{\epsilon}_{\text{MLE}}$ based on Eq. 5 with registered counts $N_{0}$ and $N_{1}$. We then obtained the MSE for each separation by averaging $(\hat{\epsilon}_{\text{MLE}}-\epsilon)^{2}$ over 50 bootstrapped experiments with the same $N$ from the total dataset of detector clicks. The uncertainty in the MSE was calculated by repeating this procedure 10 times. The detailed procedure of data collection and analysis is described in Methods.
The estimation results, including estimator standard deviations, for a total of $N=10\times 10^{3}$ detected counts are shown in Fig. 2 (the results for $2\times 10^{3}$ and $100\times 10^{3}$ counts are provided in the Supplementary Materials). The yellow markers represent the raw estimators $4\sqrt{N_{1}/N_{0}}$, which exhibit a systematic bias primarily due to the experimental mode crosstalk arising from the memory storage process and control field leakage. While these biased estimators possess low variances, their systematic bias compromises their accuracy. In contrast, the MLE estimators exhibit much lower bias compared to the raw estimator. For the smallest separations, such as $\epsilon=0$ and $0.05$, the MLE shows a non-zero bias with increased variances. The relatively larger variances (reflected by the error bars) for these small-separation MLE estimates, compared to the raw estimators, clearly demonstrate the trade-off between estimator variance (precision) and bias (accuracy), as elucidated in Eq. 6. Importantly, despite being above the ideal quantum bound, all MLE error bars consistently lie within the DI bound, providing a clear demonstration of enhanced precision relative to DI methods. To validate the platform’s unique functionality of on-demand storage, retrieval and coherent mode conversion, we also performed experiments and demonstrated super-resolving estimation with storage durations from 150 ns to 250 ns and using HG0 and HG1 retrieval modes. The experimental results are given in the Supplementary Materials.
Figure 3: Estimation performance under different photon budgets. (a) Experimental MSE times photon number $N$. The plot shows $\text{MSE}\times N$ for total detected photon counts of $2\times 10^{3}$ (orange triangles), $10\times 10^{3}$ (green squares), and $100\times 10^{3}$ (blue circles). MSE error bars are determined from 10 bootstrapped results. The solid red curve represents the theoretical CRLB for mode filtering using our Raman quantum memory, incorporating measured mode crosstalk. As the photon number increases, the MSE approaches the CRLB line. For comparison, the quantum CRLB (blue dash-dotted line) and the direct intensity detection CRLB (gray dashed line) are also shown. (b) Parameter-to-error ratio for different total detection counts. Solid lines represent theoretical PER values calculated using Eq. 7, derived from theoretical MSE values that incorporate the measured crosstalk parameters. The corresponding symbols (circles, squares, and triangles) indicate PER values computed from our experimentally measured MSE. The sensitivity, or smallest resolvable separation, is defined as the point where the PER exceeds 0 dB. Higher detection counts consistently lead to improved PER and enable the resolution of smaller frequency separations. Specifically, for $N=100\times 10^{3}$ photons, our system achieves a smallest resolvable separation of $\epsilon=0.05$.
@@
The residual bias observed for near-zero separations in MLE is directly attributed to the non-negativity constraint imposed on the parameter $\epsilon$, as we analyzed in Section A. For the limited photon budgets in our experiments ($2\times 10^{3}$ to $100\times 10^{3}$), this estimator bias is unavoidable for small separations and is dependent on the number of detected photons. Figure 3(a) illustrates the measured MSE and the CRLB, both scaled by $N$ for comparison across different total detection counts. For separations $\epsilon<0.5$, the MSE is significantly lower than the CRLB for the DI method. The experimentally measured MSE quantifies the total estimation error for a given photon budget, including contributions from both the bias and the variance of the estimator. At low photon numbers, such as $N$ = $2\times 10^{3}$, the MSE falls below the CRLB for small separations $\epsilon<0.15$ — a consequence of the estimator bias. As the photon count increases to $100\times 10^{3}$, the MSE approaches the CRLB as the bias decreases. This observation is consistent with our earlier discussion of MLE bias at small separations. These results show that for moderate photon budgets, the MLE exhibits a non-zero bias for small separations. However, the contribution of the bias to the MSE becomes negligible with increasing photon counts, allowing the estimator to perform reliable estimation at higher photon numbers.
@@ II.4 Sensitivity and precision enhancement benchmarking
Figure 4: Precision enhancement benchmarking. (a) Precision enhancement as a function of spectral line separation $\epsilon$. The red curve shows the theoretical precision enhancement, calculated as the ratio of the mode filtering CRLB to the direct intensity measurement CRLB. The symbols represent experimentally measured precision enhancements based on the MSE values obtained at various total detection counts: $N=2\times 10^{3}$ (orange triangles), $10\times 10^{3}$ (green squares), and $100\times 10^{3}$ (blue circles). The average of data points lie slightly below the theoretical curve due to additional MSE arising from experimental fluctuations. The superresolution parameter $\mathfrak{s}$ is defined as the asymptotic limit of the red curve as $\epsilon\rightarrow 0$. (b) Comparison of superesolution parameters for various schemes. This panel benchmarks the performance of our system against other super-resolution techniques and DI methods. QPG: quantum pulse gate [40]; PuDTAI: pulsedivision time-axis-inversion [43]; SUSI: super-resolution via spectral inversion [45]; QMTI: quantum-memory temporal imaging [43]; FT: Fourier transform spectrometer (Bruker IFS 125HR); Grating: grating spectrometer of grating length 1 cm and grating period of 1200 mm.
@@
To benchmark the sensitivity of our measurement scheme, we examine the minimum resolvable separation based on the PER ratio, as defined in Eq. 7. Figure 3(b) shows the PER as a function of separation for different total detection counts. Our experimental results demonstrate a PER$=4.4\pm 0.5$ dB for $\epsilon=0.05$ with $100\times 10^{3}$ photon counts. This signifies that our system can resolve spectral line separations as small as 265 kHz for spectral lines with a linewidth of 5.30 MHz. In contrast, for lower photon numbers ($N=2\times 10^{3}$ and $10\times 10^{3}$), the PER remains below 0 dB at $\epsilon=0.05$, implying that the frequency separation cannot be reliably distinguished from a zero separation at these photon levels. The solid lines in Fig. 3(b) represent the theoretical PER values, derived from theoretical MSE simulations, which show good agreement with the experimentally measured PER.
@@
In addition to sensitivity, we evaluate the precision enhancement enabled by our memory-based measurement apparatus with mode filtering, benchmarking its performance against DI methods. We quantify the enhancement as the ratio of the FI for the two methods in the limit of vanishing separation $\epsilon\to 0$. This limiting value corresponds to the superresolution parameter, $\mathfrak{s}$, which is introduced by Mazelanik et al. [43] as
$\mathfrak{s}=\lim\limits_{\epsilon\rightarrow 0}(\mathcal{F}/\mathcal{F}_{\text{DI}}).$
@@
This parameter defines the ultimate limit of precision enhancement achievable with an infinite photon budget, as a truly unbiased estimator is only realized when $N\to\infty$. This parameter can be used to assess and compare the performance of various super-resolving measurement apparatuses, with its value primarily determined by mode crosstalk. The red curve in Fig. 4(a) illustrates how the theoretical precision enhancement, represented by $\mathcal{F}/\mathcal{F}_{\text{DI}}$, varies with frequency separation for our platform. In the limit $\epsilon\rightarrow 0$, the ratio asymptotically approaches a value of approximately 37. Using this definition, our system achieves the highest reported superesolution parameter value for frequency separation estimation to date [40, 43, 45], as shown in the comparative graph in Fig. 4(b). For comparison, we also show the performance of DI methods, including quantum-memory temporal imaging [43], a Fourier transform spectrometer (Bruker IFS 125HR) and a grating spectrometer; all of these fall below $\mathfrak{s}=1$. We should also highlight that while the super-resolving methods presented span various bandwidth regimes, our Raman memory uniquely operates in the MHz to GHz range, a bandwidth dictated by the hyperfine splitting between the initial and storage states and the memory storage time.
@@
It is important to note, however, that this theoretical precision enhancement $\mathfrak{s}$ in the $\epsilon\rightarrow 0$ limit is experimentally unattainable, as achieving unbiased estimation in this regime requires an infinite number of detected photons. In practice, when the estimator exhibits negligible bias, the enhancement can be estimated using the ratio of the measured mode filtering MSE to the CRLB of DI methods. In Fig. 4(a), we plot these ratios as data points at different total photon counts. These data points align well with the tendency of the theoretical prediction, though subject to experimental imperfections and fluctuations. However, at very small separations, the estimator bias can cause the measured MSE to fall below the CRLB, leading to calculated enhancement ratios that may exceed the theoretical values (red curve in Fig. 4(a)). In this region, we expect both the mode filtering method and the DI detection method exhibit estimator biases, making the determination of realistic precision enhancement more complex [21]; a detailed analysis will be explored in future work. Our experimental measurements demonstrate a practical $(34\pm 4)$-fold improvement at $\epsilon=0.05$ and a $(28\pm 6)$-fold improvement at $\epsilon=0.1$ using $100\times 10^{3}$ detected photons, confirming the feasibility and effectiveness of the method. Furthermore, we expect DI measurements, which inherently have larger estimator variances from the FI analysis, require higher number of photons to achieve nearly unbiased estimation compared to the optimal mode filtering. The mode filtering method, with its lower variance, is especially valuable for photon-limited applications where conventional techniques fail to achieve accurate estimation.
@@ III Discussion
The experimental results demonstrate a clear improvement in the precision of estimating spectral line separation. Our system achieves a theoretical 37-fold precision enhancement in the limit $\epsilon\rightarrow 0$, and a practical $(34\pm 4)$-fold enhancement at $\epsilon=0.05$, relative to DI method for sub-linewidth separations. Our system exhibits sensitivity sufficient to resolve frequency separations as small as 1/20 of the individual signal linewidth. Among TF superresolution techniques [40, 43, 45], our Raman memory operates over a unique bandwidth, from MHz to GHz, where no classical strategy is available. Going beyond resolving two spectral lines, our platform’s fully programmable temporal mode filtering also supports the capability of resolving more complex spectral features. The user-defined mode filtering is pivotal for multi-parameter estimation, where the optimal measurement basis can be complicated depending on the task [57, 58, 59, 60]. Such optimal measurement bases are ideally realized through TF mode sorting onto multiple temporal modes [42], which could be implemented using multiple cascaded memories or a loop architecture with a single memory.
@@
As shown in Eq. 4, achieving even higher precision in our system hinges on reducing mode crosstalk, which primarily stems from the memory storage process. Crucially, there is a trade-off between mode crosstalk and storage efficiency, which arises from the low-coupling requirement for the memory’s single-mode behavior (see the Supplementary Materials for simulation results). To address this limitation, our recently developed efficiency enhancement via light-matter interference (EEVI) protocol [61] offers a path to realizing high-efficiency storage and retrieval while preserving single-modeness. Similarly, cavity-enhanced Raman memories [62, 63] can offer similar efficiencies, though their bandwidth is constrained by cavity finesse. Furthermore, employing optimal control techniques [64, 65] to tailor control pulse shapes can maximize storage efficiency for the target mode while simultaneously reducing it for orthogonal modes, thereby minimizing crosstalk.
@@
Despite the importance of precision enhancement, the practical advantage of the super-resolving measurement, quantified by FI per input photon, is significantly impacted by photon loss of the setup. To date, all experimental platforms for TF super-resolution suffer from low end-to-end efficiency. In our current setup, we measure an overall efficiency of approximately 0.3%, despite achieving an internal total efficiency of $14.6\pm 0.1\%$. This disparity primarily arises from the stringent filtering needed to suppress the co-propagating control field, which is detuned by just 9.2 GHz from the signal. To achieve an extinction ratio of $\sim 4\times 10^{7}$, we used a calcite beam displacer, a Glan-Taylor polarizer, and three double-passed Fabry–Pérot etalons, each contributing to the overall photon loss. Notably, only changes in the internal efficiency alter mode crosstalk. In contrast to Raman memories, ladder-type quantum memories, such as fast-ladder memory [66] or off-resonant cascaded absorption memory (ORCA) [67, 68], exploit a higher-lying excited state and widely separated, counter-propagating signal and control fields. This configuration, allowing efficient suppression of control leakage without complex filtering, has demonstrated noise per signal photon as low as $10^{-5}$ [67] with end-to-end efficiencies up to 35% [66]. Integrating our high-fidelity mode filtering with these architectures holds the potential for realizing unconditional TF super-resolution.
@@
In conclusion, we have demonstrated a frequency superresolution protocol based on mode-selective Raman quantum memory. The technology developed here holds promise for diverse applications, including high-precision clock synchronization [69], spacetime positioning [70, 71], photon dose-limited metrological tasks [7], and processing of TF encoded quantum information [6, 28, 27, 29]. Future works will focus on further improving precision through the implementation of the EEVI protocol and control optimization techniques. We will also explore the potential of ORCA memories to achieve unconditional super-resolving measurements. Furthermore, we aim to demonstrate the platform’s capabilities for multi-parameter estimation and its integration into quantum sensor networks. The unique combination of high-fidelity mode filtering, on-demand storage and mode conversion capabilities positions our platform as a multifunctional tool for advanced TF metrology, paving the way for integrated quantum sensors tailored to photon-limited applications and distributed sensing networks.

### v2

@@ II.2 Mode-selective Raman quantum memory
In the low-coupling regime (i.e., low storage efficiencies), the memory operates in a single-mode fashion: the stored temporal mode is directly determined by the temporal profile of the control pulse [55] (see Supplementary Information for details). Fig. 1(B) illustrates two representative cases: when both signal and control are in the $\mathrm{HG}_{0}$ mode (label 1), the signal is efficiently stored and retrieved; conversely, when the signal is in $\mathrm{HG}_{0}$ but the control is in $\mathrm{HG}_{1}$ (label 2), storage and retrieval are strongly suppressed. Experimental results for these scenarios are shown in Fig. 1(C). To quantify the mode selectivity, we measured the total efficiencies for all combinations of signal and control pulses prepared in HG0 and HG1 modes. Fig. 1(D) shows the measured mode crosstalk matrix with high mode selectivity: $(99.61\pm 0.48)\%$ when storing $\mathrm{HG}_{0}$ signals with $\mathrm{HG}_{0}$ control, and $(0.39\pm 0.02)\%$ with $\mathrm{HG}_{1}$ control. These results were obtained using read-in and read-out control pulses with energies of $125\pm 1$ pJ, yielding total efficiencies of approximately 11% along the diagonal terms. At higher coupling strengths (e.g., with stronger control pulses), this single-mode behavior degrades due to the increased storage of temporal modes orthogonal to that of the intended control field. This highlights an inherent trade-off between mode selectivity and overall memory efficiency. By achieving low crosstalk and high mode selectivity in the low-coupling regime, our Raman memory provides a robust and effective platform for the super-resolution task.
@@ II.3 Experimental superresolution with mode-selective quantum memory
Figure 2: Experimental estimation results. This graph compares the maximum likelihood estimators (red markers) with raw estimators (yellow markers), both derived from $10\times 10^{3}$ total detected photons. The dashed black line indicates the ground truth separation. The shaded regions illustrate the theoretically predicted standard deviations from the quantum CRLB (blue) and the direct intensity measurement CRLB (gray).
@@
To perform estimation experiments using the mode-selective Raman quantum memory described above, we first prepared the signal as an incoherent mixture of two Gaussian spectral lines, each with a linewidth of $\sigma=5.30$ MHz, and varied their normalized separation over 21 values from 0 to $1$ in steps of $0.05$ (see Methods). For each separation, we performed two measurements by storing the signal with either an $\mathrm{HG}_{0}$ or an $\mathrm{HG}_{1}$ control read-in pulse. In both cases, the stored signals were then retrieved after 200 ns using an $\mathrm{HG}_{1}$ pulse and detected with SNSPDs. At $\epsilon=0$, the storage efficiency and total internal efficiency for the $\text{HG}_{0}$ control read-in were measured to be $(39.5\pm 0.6)\%$ and $(14.6\pm 0.1)\%$ respectively, using control pulses with energies of $130\pm 1$ pJ. For each separation, we recorded approximately $N\approx 2\times 10^{5}$ detector clicks in total.
To estimate the frequency separation $\epsilon$, we first calibrated the experimental system with $1.6\times 10^{5}$ detected counts per separation. By fitting the experimental model to this calibration data, we determined the measured crosstalk from $\text{HG}_{0}$ to $\text{HG}_{1}$ in the estimation experiment to be 0.34%. It is noteworthy that this is a separate experiment from the mode filtering demonstration in the previous section, accounting for the slight difference in the reported crosstalk. With the calibrated model, in each run we performed MLE to compute the estimate $\hat{\epsilon}_{\text{MLE}}$ based on Eq. 3 with registered counts $(N_{0},N_{1})$. We then obtained the MSE for each separation by averaging $(\hat{\epsilon}_{\text{MLE}}-\epsilon)^{2}$ over 50 bootstrapped experiments with the same $N$ drawn from the full click dataset. The uncertainty in the MSE was evaluated by repeating this procedure 10 times.
Figure 2 shows the estimation results with estimator standard deviations for $N=10\times 10^{3}$ detected counts. The complete estimation results for various detection counts, detailing the contributions from estimator bias and variance, are presented in the Supplementary Information. While the raw estimators $4\sqrt{N_{1}/N_{0}}$ have low variance, they suffer from a systematic bias caused by mode crosstalk and control field leakage. In contrast, the MLE exhibits substantially reduced bias. For the smallest separations ($\epsilon=0-0.2$), the MLE estimators display lower bias but higher variance, reflecting the trade-off between estimator variance (precision) and bias (accuracy) described in Eq. 4. Crucially, all MLE error bars are smaller than the DI bound, demonstrating enhanced precision over DI detection. To validate the platform’s unique functionality of on-demand storage, retrieval and coherent mode conversion, we also performed super-resolving estimation with storage times from $150$ ns to $250$ ns and with $\mathrm{HG}_{0}$ and $\mathrm{HG}_{1}$ retrieval modes. The results are included in the Supplementary Information.
@@
The residual MLE bias near $\epsilon=0$ is mainly attributed to the non-negativity constraint on $\epsilon$. This estimator bias is unavoidable under limited photon budgets and depends on the number of detected photons. Fig. 3(A) plots $\mathrm{MSE}\times N$ together with the corresponding CRLBs for comparison across different detected counts. For $\epsilon<0.5$, the measured MSE is generally smaller than the DI CRLB, indicating improved precision. However, at low photon numbers (e.g. $N=2\times 10^{3}$), the MSE falls below the mode-filtering CRLB for $\epsilon<0.15$ due to estimation bias. As the photon count increases to $100\times 10^{3}$, the MSE approaches the CRLB as the bias decreases, which is consistent with our earlier discussion. Thus, for limited photon budgets the MLE incurs a small bias at small separations, but its contribution to the MSE becomes negligible as $N$ increases, allowing reliable, high-precision estimation at larger photon numbers.
Figure 3: Estimation performance under different photon budgets. (A) MSE scaled by photon number NN. Experimental results for $N=2\times 10^{3}$ (orange triangles), $10\times 10^{3}$ (green squares), and $100\times 10^{3}$ (blue circles) approach the theoretical CRLB of our setup (solid red line) as NN increases. The quantum CRLB (blue dash-dotted line) and the direct intensity detection CRLB (gray dashed line) are shown for reference. Error bars are from 10 bootstrapped results. (B) Parameter-to-error ratio for different detection counts. Experimental data (markers) are plotted alongside their corresponding theoretical values (solid lines). The dashed line at 0 dB marks the sensitivity threshold.
@@ II.4 Sensitivity and precision enhancement benchmarking
To benchmark the sensitivity of our measurement scheme, we examine the minimum resolvable separation based on the PER ratio, as defined in Eq. 5. Fig. 3(B) shows the PER as a function of separation for different total detection counts. With $100\times 10^{3}$ detected photons, we achieve $\mathrm{PER}=4.4\pm 0.5$ dB at $\epsilon=0.05$, corresponding to resolving spectral lines separated by just $265$ kHz for a linewidth of $5.30$ MHz. In contrast, for lower photon numbers ($N=2\times 10^{3}$ and $10\times 10^{3}$), $\mathrm{PER}<0$ dB at $\epsilon=0.05$, indicating that such small separations cannot be distinguished from zero separation at these photon levels. The theoretical PER values (solid lines in Fig. 3(B)) show good agreement with the experimental results.
@@
In addition to sensitivity, we evaluate the precision enhancement enabled by our memory-based mode filtering scheme against DI methods. We quantify the enhancement via the superresolution parameter $\mathfrak{s}$ introduced by Mazelanik et al. [43], which is defined as the ratio of the FI for the two methods in the limit of vanishing separation $\epsilon\to 0$
$\mathfrak{s}=\lim\limits_{\epsilon\rightarrow 0}(\mathcal{F}/\mathcal{F}_{\text{DI}}).$ (6)
@@
This parameter defines the precision enhancement achievable with an infinite photon budget, as a truly unbiased estimator is only realizable when $N\to\infty$. This parameter can be used to assess and compare the performance of various super-resolving measurement schemes, with its value primarily determined by mode crosstalk. Figure 4(A) shows how the theoretical enhancement $\mathcal{F}/\mathcal{F}_{\mathrm{DI}}$ (red curve) varies with $\epsilon$. In the limit $\epsilon\rightarrow 0$, the ratio asymptotically approaches $\sim 37$ for our platform. This value sets a new benchmark for frequency-separation estimation, surpassing previously reported results [40, 43, 45], as shown in the comparative graph in Fig. 4(B). For comparison, we also show the performance of DI methods — including quantum-memory temporal imaging [43], a Fourier transform spectrometer and grating spectrometer - all of which exhibit $\mathfrak{s}<1$. Among techniques spanning various bandwidth regimes, our Raman memory uniquely operates in the MHz to GHz range, set by the hyperfine splitting and the memory storage time.
@@
It should be noted that under finite statistics, the experimental precision enhancement may deviate from the theoretical values due to estimator bias. In practice, when the estimator exhibits negligible bias, the enhancement can be approximated using the ratio of the measured mode filtering MSE to the CRLB of DI methods. Shown as data points in Fig. 4(A), these ratios align well with the trend of the theoretical prediction, subject to experimental imperfections and fluctuations. However, at very small separations, the estimator bias can cause the measured MSE to fall below the CRLB, resulting in calculated ratios that exceed the theoretical values. In this region, we expect both the mode filtering method and the DI detection method exhibit estimator biases, making the determination of realistic precision enhancement more complex [21]; a detailed analysis will be explored in future work. Our experiment demonstrates a practical $(34\pm 4)$-fold improvement at $\epsilon=0.05$ and a $(28\pm 6)$-fold improvement at $\epsilon=0.1$ using $100\times 10^{3}$ detected photons, confirming the feasibility and effectiveness of our approach. Furthermore, we expect DI measurements, which inherently have larger estimator variances from the FI analysis, require higher number of photons to achieve nearly unbiased estimation, whereas the optimal mode filtering maintains lower variance even in photon-limited applications.
Figure 4: Precision enhancement benchmarking. (A) Precision enhancement as a function of spectral line separation $\epsilon$. The theoretical precision enhancement (red curve) is plotted with all experimental results for various photon counts: $N=2\times 10^{3}$ (orange triangles), $10\times 10^{3}$ (green squares), and $100\times 10^{3}$ (blue circles). (B) Comparison of superesolution parameters for various schemes. This panel benchmarks the performance of our system against other super-resolution techniques and DI methods. QPG: quantum pulse gate [40]; PuDTAI: pulsedivision time-axis-inversion [43]; SUSI: super-resolution via spectral inversion [45]; QMTI: quantum-memory temporal imaging [43]; FT: Fourier transform spectrometer (Bruker IFS 125HR); Grating: grating spectrometer of grating length 1 cm and grating period of 1200 mm.
@@ III Discussion
We have demonstrated a clear improvement in the precision of spectral line separation estimation. Our system achieves a practical sensitivity sufficient to resolve separations as small as 1/20 of the individual signal linewidth, providing a $(34\pm 4)$-fold precision enhancement over direct-intensity methods at this separation. Among TF superresolution techniques [40, 43, 45], our Raman memory operates over a unique bandwidth, from MHz to GHz, where no classical strategy is available. Going beyond resolving two spectral lines, our platform’s fully programmable temporal mode filtering also opens the door to resolving more complex spectral features. The user-defined mode filtering is pivotal for multi-parameter estimation, where the optimal measurement basis can be complicated depending on the task [56, 57, 58, 59]. Such optimal measurement bases are ideally realized through TF mode sorting onto multiple temporal modes [42], which could be implemented using multiple cascaded memories or a loop architecture with a single memory.
@@
As indicated in Eq. 2, achieving even higher precision hinges on reducing mode crosstalk, which primarily stems from the memory storage process. Crucially, there is a trade-off between mode crosstalk and storage efficiency, which arises from the low-coupling requirement for the memory’s single-mode behavior (see Supplementary Information for simulation results). To address this limitation, our recently developed efficiency enhancement via light-matter interference (EEVI) protocol [60] offers a path to realizing high-efficiency storage and retrieval while preserving single-modeness. Cavity-enhanced Raman memories provide a similar solution [61, 62], though their bandwidth is constrained by cavity finesse. Alternatively, optimal control techniques can tailor pulse shapes to maximize storage efficiency for the target mode while suppressing orthogonal modes, thereby minimizing crosstalk [63, 64]. These strategies offer promising pathways toward overcoming the trade-off between storage efficiency and crosstalk.
@@
Despite the importance of precision enhancement, the practical advantage of the super-resolving measurement, quantified by FI per input photon, is greatly impacted by the photon loss of the setup. To date, all experimental platforms for TF super-resolution suffer from low end-to-end efficiency. In our current setup, we measure an overall efficiency of approximately 0.3%, despite achieving an internal total efficiency of $(14.6\pm 0.1)\%$. This disparity is caused by the stringent filtering needed to suppress the co-propagating control field, detuned by just 9.2 GHz. Achieving a high extinction ratio of $\sim 4\times 10^{7}$ requires a series of lossy optics, including a calcite beam displacer, a Glan-Taylor polarizer, and three double-passed Fabry–Pérot etalons. In contrast to Raman memories, ladder-type quantum memories, such as fast-ladder memory [65] or off-resonant cascaded absorption memory (ORCA) [66, 67], exploit a higher-lying excited state and widely separated, counter-propagating signal and control fields. This configuration, allowing efficient suppression of control leakage without complex filtering, has demonstrated noise per signal photon as low as $10^{-5}$ [66] with end-to-end efficiencies up to 35% [65]. Integrating our high-fidelity mode filtering with these architectures holds the potential for realizing unconditional TF super-resolution.
@@
In conclusion, we have demonstrated a frequency superresolution protocol using a mode-selective Raman quantum memory. The technology developed here holds promise for diverse applications, including high-precision clock synchronization [68], spacetime positioning [69, 70], photon dose-limited metrological tasks [7], and processing of TF encoded quantum information [6, 28, 27, 29]. Future works will focus on further improving precision through the implementation of the EEVI protocol and control optimization techniques. We will also explore the potential of ORCA memories to achieve unconditional super-resolving measurements. Furthermore, we hope to demonstrate the platform’s capabilities for multi-parameter estimation and its integration into quantum sensor networks. Ultimately, the unique combination of high-fidelity mode filtering, on-demand storage and mode conversion capabilities positions our platform as a multifunctional tool for advanced TF metrology, paving the way for integrated quantum sensors tailored to photon-limited applications and distributed sensing networks.